OpenStudy (julierosehead):

A. If Mia rides her bike at 14 miles per hour for 1.5 hours and Chloe drives her car at 35 miles per hour for 45 minutes, who travels farther? B. Would the person who has gone the shorter distance be able to catch up within 30 minutes if the other stops? thank you! :)

OpenStudy (hayisforhorses):

Make them both into linear equations. Mia is y=14x and Chloe is y=35x. Mia went 1.5 hours and Chloe went 0.75 miles. 14*1.5=21 and 35*0.75=26.25. So far, Chloe has gone farther. However, in 30 minutes Mia will go another 7 miles - taking her to 28 miles. This is larger than Chloe's 26.25.

OpenStudy (julierosehead):

what equation did you use for the b part?

OpenStudy (hayisforhorses):

I simply used 2 whole hours into Mia's old equation. She went 1.5 hours, then another 0.5 hour, so 14*2=28. Chloe still had the same distance from part A.

OpenStudy (julierosehead):

thank you so much :)